# All-optical compact setup for generation of collimated multi-MeV proton beams with a ”snail” target

### v1

@@ II Magnetic field structure and parameters
The current geometry providing the magnetic field structure is shown in Fig. 1, panels (b,c), with arrows. The discharge current is marked with the light green arrow, the current of deflected laser-accelerated electrons is shown with the light-blue arrow, and the red arrow indicates the laser beam propagation. The target profile is shown with the dark gray mask. Both the uniform and the coaxial magnetic field profiles are shown in the figure. The electric field, which is also excited in the system as the target is being charged positively under the action of the intense driver Ehret et al. (2022), is also taken into account, though its effect in the considered situation is weak.
@@
The propagation of high-energy proton beams through the considered electromagnetic structures was modeled using a test-particle approach. The simulations were performed using a self-developed code written in the high-level programming language Python Van Rossum and Drake (2009), complemented by the NumPy library Harris et al. (2020) and an open-source JIT-compiler Numba Lam, Pitrou, and Seibert (2015) transforming a regular Python code into a fast machine code and allowing parallelization for multiple CPUs. A beam of protons was sent through the cavity of the "snail" target along the axis ($z$-axis in Fig. 1, (b,c)). For the propagating particles in the given fields, relativistic equations of motion with the Lorentz force were solved numerically for each particle in the beam using the Boris integration scheme Boris (1970). The fields were assumed to be stationary during the time required for a multi-MeV proton to cross the target region ($100$ MeV protons cross $\simeq 100$ $\upmu$m-long high field region in less than $1$ ps, while the magnetic field changes much slower, on the temporal scale of a few hundreds of ps Ehret et al. (2022); Bukharskii and Korneev (2023)). The interaction between different particles in the proton beam was neglected as space-charge effects are typically low for TNSA-produced proton beams Santos et al. (2018). Proton scattering and energy losses due to collisions with plasma formed in the target cavity were also assumed to be negligible for the considered projectile energies exceeding $10$ MeV scales, taking into account small plasma cloud spatial scales of $\sim 100$ $\upmu$m and characteristic density values of $10^{19}$..$10^{21}$ cm-3. The magnetic fields in the target region were calculated on a 3D dimensional grid with 2 $\upmu$m resolution under the assumption that they are produced by the fixed electric currents defined by the contours in Fig. 1, (b,c). The current value was $\approx 7\cdot 10^{5}$ A, which corresponds to the $10^{5}$ T scale magnetic fields obtained in 2D PIC simulations in Bukharskii and Korneev (2023). The electric fields were determined under the assumption that the target is a perfect conductor charged to a certain potential. The latter was set to $200$ kV, similar in scale to the value retrieved from the experimental data in Ehret et al. (2022). Although, it should be noted that for the considered $\simeq 100$ MeV protons electric fields corresponding to potentials of a few hundred kilovolts have a little effect on trajectories of fast protons – further modeling of proton beam propagation showed the resulting beam profile remained almost unchanged if the electric potential was varied in $\left(0-1000\right)$ kV range. Particles were originated from a point source placed $0.5$ mm away from the "snail" center, see the related discussion in Section LABEL:sec4. Their velocities were defined according to their kinetic energy, and their initial angles were set so that the angular distribution of the beam presented a Gaussian profile with the full width at half maximum (FWHM) of $20^{\circ}$, which corresponds to the half opening angle $\theta/2=10^{\circ}$. The beam was directed along the "snail" axis and registered at the "detector" plane placed $50$ mm away from the the target plane. A total of $10^{6}$ particles were used in the simulation to accumulate sufficient statistics in the detector plane.
@@ III Proton beam collimation in the electromagnetic fields of the "snail" target
IV Conclusion
@@
The proposed method can be used at modern and perspective petawatt and multi-petawatt laser facilities (e.g. Danson et al. (2019); Kostyukov et al. (2023)) as a part of an integrated all-optical acceleration and guiding setup. In such a scheme, e.g. the TNSA-produced protons created by one laser pulse may pass the optically-driven electromagnetic structure of the considered "snail"-type target. Such a source of fast protons is expected to have an exceptional brightness, low divergence and high energy of accelerated particles, and it would be of interest for fundamental studies and prospective applications.
V Acknowledgements

### v2

@@ II Magnetic field structure and parameters
The current geometry providing the magnetic field structure is shown in Fig. 1, panels (b,c), with arrows. The discharge current is marked with the light green arrow, the current of deflected laser-accelerated electrons is shown with the light-blue arrow, and the red arrow indicates the laser beam propagation. The target profile is shown with the dark gray mask. Both the uniform and the coaxial magnetic field profiles are shown in the figure. The electric field is of course also excited in the system as the target is being charged positively under the action of the intense driver Ehret et al. (2022), though its effect in the considered situation is weak, see the discussion below.
@@
The propagation of high-energy proton beams through the considered electromagnetic structures was modeled using a test-particle approach. The simulations were performed using a self-developed code written in the high-level programming language Python Van Rossum and Drake (2009), complemented by the NumPy library Harris et al. (2020) and an open-source JIT-compiler Numba Lam, Pitrou, and Seibert (2015) transforming a regular Python code into a fast machine code and allowing parallelization for multiple CPUs. A beam of protons was sent through the cavity of the "snail" target along the axis ($z$-axis in Fig. 1, (b,c)). For the propagating particles in the given fields, relativistic equations of motion with the Lorentz force were solved numerically for each particle in the beam using the Boris integration scheme Boris (1970). The fields were assumed to be stationary during the time required for a multi-MeV proton to cross the target region ($100$ MeV protons cross $\simeq 100$ $\upmu$m-long high field region in less than $1$ ps, while the magnetic field changes much slower, on the temporal scale of a few hundreds of ps Ehret et al. (2022); Bukharskii and Korneev (2023)). The interaction between different particles in the proton beam was neglected as space-charge effects are typically low for TNSA-produced proton beams Santos et al. (2018). Proton scattering and energy losses due to collisions with plasma formed in the target cavity were also assumed to be negligible for the considered projectile energies exceeding $10$ MeV scales, taking into account small plasma cloud spatial scales of $\sim 100$ $\upmu$m and characteristic density values of $10^{19}$..$10^{21}$ cm-3. The magnetic fields in the target region were calculated on a 3D dimensional grid with 2 $\upmu$m resolution under the assumption that they are produced by the fixed electric currents defined by the contours in Fig. 1, (b,c). For this purpose, each of the assumed current paths, i.e. one in the case of the uniform B-field profile (light green in Fig. 1(b)) or two in the case of the coaxial B-field profile (light green and light blue in Fig. 1(c)), were divided into $N_{I}=100$ infinitely thin linear segments of approximately equal lengths. The resulting magnetic field at position $\mathbf{r}$ was found as a sum of contributions from all the $N_{I}$ segments according to the Biot-Savart law:
$\mathbf{B}\left(\mathbf{r}\right)=\frac{\mu_{0}}{4\pi}\sum_{i=1}^{N_{I}}\frac{I\,d\mathbf{l}_{i}\times\mathbf{r}_{i}^{\prime}}{\lvert\mathbf{r}_{i}^{\prime}\rvert^{3}},$
where $I$ is the current value, $d\mathbf{l}_{i}$ is the vector drawn from the start to the end of the $i$-th linear segment, $\mathbf{r}_{i}^{\prime}=\mathbf{r}-\mathbf{r}_{i}$ is the vector from the position $\mathbf{r}_{i}$ of the $i$-th linear segment to the point $\mathbf{r}$ where the field is calculated, and $\mu_{0}$ is the magnetic constant. The current value was $\approx 7\cdot 10^{5}$ A, which corresponds to the $10^{5}$ T scale magnetic fields obtained in 2D PIC simulations in Bukharskii and Korneev (2023). The electric fields were determined under the assumption that the target is a perfect conductor charged to a certain potential. The latter was set to $200$ kV, similar in scale to the value retrieved from the experimental data in Ehret et al. (2022). Although, it should be noted that for the considered $\simeq 100$ MeV protons electric fields corresponding to potentials of a few hundred kilovolts have a little effect on trajectories of fast protons – further modeling of proton beam propagation showed the resulting beam profile remained almost unchanged if the electric potential was varied in $\left(0-1000\right)$ kV range. Particles were originated from a point source placed $0.5$ mm away from the "snail" center, see the related discussion in Section IV. Their velocities were defined according to their kinetic energy, and their initial angles were set so that the angular distribution of the beam presented a Gaussian profile with the full width at half maximum (FWHM) of $20^{\circ}$, which corresponds to the half opening angle $\theta/2=10^{\circ}$. The beam was directed along the "snail" axis and registered at the "detector" plane placed $50$ mm away from the the target plane. A total of $10^{6}$ particles were used in the simulation to accumulate sufficient statistics in the detector plane.
@@ III Proton beam collimation in the electromagnetic fields of the "snail" target
The obtained collimation properties of the target may be rather qualitatively described in terms of a magnetic lens approach. Following Reiser (2008), in the paraxial approximation, the focal length $f$ for a solenoidal magnetic field lens with the peak magnetic field $B_{0}$ and the effective field length
[…]
is
$f=\frac{l}{\Phi\sin\Phi},$ (1)
[…]
@@
IV Discussion
The calculations described in Sec. III were performed under the assumption that the magnetic field induced around the target remains static and thus does not excite vortex electric field. To justify the validity of such assumption for the considered problem, an additional set of simulations was performed to study the effect of the inductive electric field on particle collimation. This field was calculated as
$\mathbf{E_{i}}=-\frac{1}{c}\frac{\partial\mathbf{A}}{\partial t},$ (2)
where $\mathbf{A}$ denotes the retarded vector potential:
$\mathbf{A}\left(\mathbf{r},t\right)=\frac{1}{c}\int_{\Omega}\frac{\mathbf{J}\left(\mathbf{r^{\prime}},t^{\prime}\right)}{\left|\mathbf{r}-\mathbf{r^{\prime}}\right|}d^{3}\mathbf{r^{\prime}}.$ (3)
The current in Eq. (3) was constant along each of the selected paths, see Fig. 1, panels (b,c), and decayed linearly in time from $\sim 7\cdot 10^{5}$ A (corresponding magnetic field is $B\sim 10^{5}$ T), down to zero. To demonstrate the effect, different decay times were considered, $\tau_{d}=100$ ps, $\tau_{d}=10$ ps and $\tau_{d}=1$ ps, where the value $\tau_{d}=100$ ps corresponds to the one retrieved from experimental data for "snail" targets in Ehret et al. (2022). Resulting beam profiles are shown in Fig. 5.
Figure 5: Results of test-particle simulations of proton beam transport through the "snail" cavity for different B-field profiles, uniform (top row) and coaxial (bottom row), with the inductive electric field caused by the time-varying magnetic field taken into account. Panels (a,d) correspond to the magnetic field decay time $\tau_{d}=100$ ps, panels (b,e) – to $\tau_{d}=10$ ps and panels (c,f) – to $\tau_{d}=1$ ps.
The inductive electric field induced by the decaying annular current is also annular in the target cavity region. Its main component is $E_{\phi}$ where $\phi$ denotes the azimuthal angle measured in the plane perpendicular to the beam propagation axis. Such electric field changes projectile’s azimuthal velocity component, which in turn causes a change of the radial component velocity component, as $\mathbf{v}_{\phi}\times\mathbf{B}$ force is directed radially inside the "snail" cavity. From the obtained distributions for the uniform B-field profile it follows that if the B-field changes on the time-scale $\sim\left(10-100\right)$ ps, the described effect is insignificant, see panels (a) and (b) in Fig. 5, where the beam profiles change only very slightly in comparison to the those presented in Sec. III, with a decrease of peak proton density at the center of the collimated beam not exceeding a few percent. For $\tau_{d}=1$ ps, however, the inductive electric field strength appears to be strong enough to cause a significant redistribution of protons and reduce the collimation efficiency. This case is shown in panel (c) of Fig. 5. The same applies to the beam profiles obtained for the coaxial B-field structure, shown in panels (d-f). As for the uniform B-field profile, the inductive electric field impairs collimation only for $\tau_{d}=1$ ps, see panel (f) in Fig. 5. Therefore, for the significant inductive electric field contribution in the considered scheme, the magnetic field has to change on $\sim 1$ ps time-scale, which is much faster than the B-field decay rate measured experimentally Ehret et al. (2022) and observed in simulations Bukharskii and Korneev (2023). Thus, in the considered range of parameters, the inductive fields do not play a significant role in the beam collimation. The static electric fields for realistic target potentials up to $1$ MV also appear to be weak to cause a significant deflection of $\simeq 100$ MeV protons, as it was verified by an auxiliary set of simulations. Thus, we can conclude that for the considered high energies of protons and strong values of the magnetic field, the collimation effect is mainly defined by the quasistationary magnetic field induced around the target.
Another important point worth discussing is the used model of the proton beam – in Sec. III a simple ideal case was considered, where a monoenergetic proton beam is emitted from a point-like source, while in reality the source has finite dimensions and some energy spread. Assuming that the proton beam is produced via the TNSA mechanism during an ultrashort laser pulse irradiation of a thin metallic foil. The imaging properties of such a source, despite relatively large transverse size of the emitting region of $\sim 100-200$ $\upmu$m, can be reproduced by assuming a much smaller source shifted by $\sim 100$ $\upmu$m from the foil surface Borghesi et al. (2004). The effective size of such a source can be as small as $\simeq 5-10$ $\upmu$Borghesi et al. (2004); Kugland et al. (2012). As discussed in Kugland et al. (2012), finite source size may lead to blurring of proton images, i.e. beam profiles, and can be neglected if the source size $d_{s}$ is much lower than the probed field size $d$: $d_{s}\ll d$. Here, assuming $d_{s}\simeq 5-10$ $\upmu$m the condition $d_{s}\ll d$ is satisfied as $d\simeq 50$ $\upmu$m, and blurring related to the finite source size should be insignificant, which was verified by an auxiliary set of simulations. There, the initial transverse proton positions were chosen randomly inside circles with different diameters, $d_{s}=0.2$ $\upmu$m, which corresponds to an almost point-like source, as well as $d_{s}=5$ $\upmu$m and $d_{s}=10$ $\upmu$m which describe finite-size sources. Results of the performed simulations for the uniform B-field profile are shown in Fig. 6. As can be seen, blurring becomes significant only for $d_{s}=10$ $\upmu$m, see Fig. 6(c). In this case, the peak proton density drops by $\approx 35$ in comparison to the point-like source case, shown in panel (a) of the same figure. For $d_{s}=5$ $\upmu$m, the peak proton density decrease due to the blurring amounts to a modest $10$. Similar results were obtained assuming the coaxial B-field profile. Thus, for the considered scheme the point source assumption remains viable for $d_{s}\lesssim 5$ $\upmu$m as the transverse source size in this case is much smaller than the collimating field transverse size, and the collimation efficiency remains almost unaffected.
Figure 6: Results of test-particle simulations of proton beam transport through the "snail" cavity for the uniform B-field, $120$ MeV proton energy and different source sizes $d_{s}$: (a) point-like source with $d_{s}=0.2$ $\upmu$m, (b,c) finite-size sources with $d_{s}=5$ $\upmu$m and $d_{s}=10$ $\upmu$m, respectively.
In terms of the energy spectrum, TNSA-produced protons are usually characterized by a wide energy spread Roth and Schollmeier (2016). Although, there have been a number of reports of quasi-monoenergetic ion beams produced via TNSA-mechanism using specifically-designed and prepared targets Hegelich et al. (2006); Schwoerer et al. (2006); Ter-Avetisyan et al. (2006) – there, the energy spread of accelerated ions did not exceed $20$. Despite relative difficulties in obtaining monoenergetic ion beams with TNSA, it should be mentioned that for practical applications involving interaction of the TNSA-produced proton beams with thick targets the target itself plays a role on an energy filter due to the presence of well-localized region of ion losses in matter, the so-called Bragg peak. In this case, only a small fraction of the whole beam energy from a certain part of the spectrum, is deposited in the target at a certain depth. This allows for the used description of the monoenergetic proton beams in Sec. III. Nevertheless, it is worth studying how the energy spread effects the collimation efficiency. The study was performed both for the uniform and coaxial B-field profiles, assuming three different values of the energy spread – $\delta E_{0}=\pm 1$, i.e. almost a monoenergetic beam, as well as $\delta E_{0}=\pm 20$ and $\delta E_{0}=\pm 50$. Results are summarized in Fig. 7, panels (a-f).
Figure 7: Results of test-particle simulations of a non-monoenergetic proton beam transport through the "snail" cavity: (a-c) $120$ MeV protons in the uniform B-field with different values of energy spread, $\delta E_{0}=1$, $\delta E_{0}=20$ and $\delta E_{0}=50$, respectively; (d-f) $110$ MeV protons in the coaxial B-field with the same values of the energy spread, $\delta E_{0}=1$, $\delta E_{0}=20$ and $\delta E_{0}=50$, respectively; (g-h) beam profiles for the uniform and coaxial B-field profiles assuming an exemplary angular-energy distribution for the TNSA-mechanism.
From the obtained distributions it follows that the decrease of the collimation efficiency is relatively modest, even for $\delta E_{0}=50$ – the collimated spot retains almost the same shape as in the case of a quasi-monoenergetic beam shown in Fig. 7(a,d), but the peak proton density at the center of this spot slightly drops. For $\delta E_{0}=20$ which has already been successfully achieved in several experiments on laser-driven ion acceleration Hegelich et al. (2006); Schwoerer et al. (2006); Ter-Avetisyan et al. (2006), this drop amounts to only $\simeq\left(6-12\right)$, see Fig. 7(b,e). For higher $\delta E_{0}$ values, the decrease of the peak proton density depends on the B-field structure – for the uniform profile, it still remains relatively small, about $12$, see Fig. 7(c), while for the coaxial profile it amounts to $\approx 30$, see Fig. 7(f).
[…]
Finally, it is also worth discussing how the beam profile changes as it propagates along its axis. Above, distributions in a single plane at a fixed distance of $50$ mm from the target plane were provided. To illustrate the beam evolution as is propagates in the longitudinal direction, transverse beam profiles were compared at different distances between the detector and the target plane – $40$ mm, $50$ mm and $60$ mm (i.e. in ±$\pm 10$ mm distance range from the plane where the profiles had been calculated previously) and additionally at $250$ mm (i.e. $5$ times farther from the target than before). The results are shown in Fig. 8.
Figure 8: Results of test-particle simulations of proton beam transport through the magnetized "snail" target: (a-c) beam profiles for $120$ MeV protons and for the uniform B-field at different distances from the target plane – $40$ mm, $50$ mm and $60$ mm, respectively; (d-f) beam profiles for $110$ MeV protons and the coaxial B-field at the same three distances from the target plane – $40$ mm, $50$ mm and $60$ mm, respectively; (g) the beam profile for $120$ MeV protons and the uniform B-field at $250$ mm distance from the target plane; (h) the beam profile for $110$ MeV protons and the coaxial B-field at $250$ mm distance from the target plane.
Examination of the subsequent transverse cuts at $\left(50\pm 10\right)$ mm both for the uniform (Fig. 8(a-c)) and the coaxial B-field profiles (Fig. 8(d-f)) indicates that the collimated beam in not completely parallel and retains some residual divergence, as was already illustrated, e.g. in panels (c) of Figs. 2 and 3. This results in a linear decrease of the central spot size with distance, so that at the $250$ mm distance it appears to be about $5$ times larger than at the $50$ mm distance, while the peak proton concentration drops as the distance squared, so that $250$ mm away from the target it is about twenty times lower than at the $50$ mm distance, see Fig. 8(a-f) and Fig. 8(g,h). This can be in part a consequence of an imperfectly symmetric field structure, i.e. asymmetry of the field distribution relative to the beam propagation axis. Yet, the imposed magnetic field has a strong positive effect on the peak proton density – in the central spot corresponding to protons that pass through the target cavity for all distances it increases by $\simeq\left(7-14\right)$ times, depending on the field structure. At $250$ mm distance from the target, the inner structure of the central spot visible in panels (a-f) becomes resolved, see panels (g,h) of Fig. 8. Interestingly, it may also have additional, "inner" caustics, which are better visible in the case of the uniform field, see panel (g). A small spot formed by collimated protons is visible close to the center of the profile in both panels (g) and (h). At a large distance it can be seen that this most collimated proton fraction slightly changes its propagation direction – the spot is shifted from $\left(0,0\right)$ position at the detector plane, by the shift is small, with the deflection angle relative to the propagation axis of $\simeq\left(0.3^{\circ}-0.4^{\circ}\right)$.
V Conclusion
@@
The collimation effect for the scheme considered here is attributed mainly to the action of the $\sim 10^{5}$ T scale magnetic field, while static and inductive electric fields for realistic parameters do not have a significant impact on the collimation efficiency. The effect of finite source size appears to be negligible, as long as it stays at least an order of magnitude lower than the size of the collimating cavity. Here, in order not to impair the collimation efficiency, the effective source size has to be $\lesssim 5$ $\upmu$m. Although, the considered system can be used to collimate TNSA-produced beams with a wide energy spectrum, it is better suited for the collimation of quasi-monoenergetic beams with energy spread not exceeding $\simeq 20$. However, collimation of beams with wide energy spectrum can be improved by fine-tuning of the B-field decay time – it can be optimized in such a way, that the field’s decay rate matches the time of flight for protons with a particular energy to provide their efficient collimation, turning the considered system into an achromatic magnetic lens. In this case, protons with different energies are assumed to approach the magnetic field region in different times, the lower the energy the greater this time, depending on the initial proton source position. For slower protons the magnetic field needed for their efficient collimation is lesser, and the suitable values may be in principle adjusted using different materials, thicknesses or structures of the targets.
It was found that although, the beam retains some residual divergence $\simeq 1^{\circ}$, probably as a consequence of its non-axisymmetric structure, the magnetic field of the snail target still greatly increases proton concentration at the center of the detector plane – a $\simeq\left(7-14\right)$ increase is observed at all distances from the target plate relative to the initial divergent beam case. The proposed method can be used at modern and perspective petawatt and multi-petawatt laser facilities (e.g. Danson et al. (2019); Kostyukov et al. (2023)) as a part of an integrated all-optical acceleration and guiding setup. In such a scheme, e.g. the TNSA-produced protons created by one laser pulse may pass the optically-driven electromagnetic structure of the considered "snail"-type target. Such a source of fast protons is expected to have an exceptional brightness, low divergence and high energy of accelerated particles, and it would be of interest for fundamental studies and prospective applications.
VI Acknowledgements